# Linking numbers of modular knots

### 02 Jan 2024 v2

@@ Abstract.
The linking numbers between modular knots and the trefoil are well understood. Indeed, É. Ghys showed in [Ghy07] that they are given by the Rademacher invariant of the corresponding conjugacy classes. The Rademacher function $\operatorname{Rad}\colon\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}$ is a homogeneous quasi-morphism which he recognised with J. Barge in [BG92] as half the primitive of the bounded Euler class. This shed light on the work [Ati87] of M. Atiyah which identified it with no less than six other important functions appearing in diverse areas of mathematics.
@@
This paper contains the main results obtained in the second part of my thesis. I would thus like to thank my thesis advisors Etienne Ghys and Patrick Popescu-Pampu for their guidance and encouragement; as well as Francis Bonahon, Louis Funar, Jean-Pierre Otal and Anne Pichon who refereed and carefully read my work. I am also grateful to Pierre Dehornoy for sharing his knowledge of modular knots. I owe Marie Dossin for helping me with the figures in tikz. Finally, I thank the referee for their relevant suggestions and questions.
@@ 0. Introduction
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ acts properly discontinuously on the hyperbolic plane $\mathbb{H}\mathbb{P}$ with quotient the modular orbifold $\mathbb{M}$, a hyperbolic surface with conical singularities $i$ $j$ of order $2$ $3$, and a cusp $\infty$. The free homotopy classes of loops in $\mathbb{M}$ correspond to the conjugacy classes of its fundamental group $\pi_{1}(\mathbb{M})=\operatorname{PSL}_{2}(\mathbb{Z})$. In particular the hyperbolic conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the closed oriented geodesics in $\mathbb{M}$, called modular geodesics. For hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ the modular geodesic $\alpha$ has length $\lambda_{A}$ equal to the logarithm of the ratio between its eigenvalues $\epsilon_{A}^{\pm 1}$, in formula:
@@
$L=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\qquad\&\qquad R=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$
and more precisely that its submonoid $\operatorname{SL}_{2}(\mathbb{N})$ of matrices with non-negative entries is freely generated by $L\&R$. This submonoid can be identified with its image $\operatorname{PSL}_{2}(\mathbb{N})\subset\operatorname{PSL}_{2}(\mathbb{Z})$. In $\operatorname{PSL}_{2}(\mathbb{Z})$, the conjugacy class of an infinite order element intersects $\operatorname{PSL}_{2}(\mathbb{N})$ along all cyclic permutations of a non-empty $L\&R$-word. The conjugacy class is primitive if and only if the cyclic word is primitive, and it is hyperbolic when the cyclic word contains both letters $L$ and $R$.
One may try to relate the geometry and topology of the master modular link with the arithmetics and combinatorics of conjugacy classes in the modular group. Our previous work [Sim22b] relates the geometry of modular geodesics (angles of intersection and lengths of ortho-geodesics) in terms of the arithmetics of conjugacy classes in the modular group (discriminants, cross-ratios between fixed points on the projective line, and their Hilbert symbols). The main results in this paper will relate the linking numbers of modular knots to the combinatorics of the corresponding cyclic words.
@@
The arithmetic & geometry of the cosines
The journey from arithmetics began with our previous work [Sim22b]. In particular for a field $\mathbb{K}\supset\mathbb{Q}$, we described when two hyperbolic elements $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ of the same discriminant $\Delta$ are conjugate in $\operatorname{PSL}_{2}(\mathbb{K})$. The obstruction is measured in terms of any intersection angle $\theta$ between their modular geodesics $[\alpha],[\beta]$ by the class of $\left(\cos\tfrac{\theta}{2}\right)^{2}\in\mathbb{K}^{\times}$ modulo the group of norms over $\mathbb{K}$ of the extension $\mathbb{K}[\sqrt{\Delta}]$. When this obstruction vanishes, the elements of $\operatorname{PSL}_{2}(\mathbb{K})$ which conjugate $A$ to $B$ are paremetrized by the points $(X,Y)\in\mathbb{K}$ of the generalised Pell-Fermat conic with equation $X^{2}-\Delta Y^{2}=\left(\cos\tfrac{\theta}{2}\right)^{2}$. Hence the geometric quantities $\left(\cos\tfrac{\theta}{2}\right)^{2}=\tfrac{1+\cos(\theta)}{2}$ given for some representatives $A,B\in\operatorname{SL}_{2}(\mathbb{R})$ whose axes intersect in $\mathbb{H}\mathbb{P}$ by
@@
$\operatorname{C}_{q}(A,B)\xrightarrow[q\to\infty]{}\operatorname{lk}(A,B)-\operatorname{lk}(A^{-1},B)=\operatorname{lk}(A,B)-\tfrac{1}{4}I(A,B)$
Hence the functions $\operatorname{L}_{q}\&\operatorname{C}_{q}$ interpolate between the geometry at $q=1$ of the arithmetic group $\operatorname{PSL}_{2}(\mathbb{Z})\subset\operatorname{PSL}_{2}(\mathbb{R})$ and the topology at $q=\infty$ of the combinatorial action $\operatorname{PSL}_{2}(\mathbb{Z})\to\operatorname{Aut}(\mathcal{T})$.
@@
A primitive hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})=\pi_{1}(\mathbb{M})$ corresponds to a primitive modular geodesic in $\alpha\subset\mathbb{M}$. It lifts to a modular knot in $\vec{\alpha}\subset\mathbb{U}$ which in turns yields a conjugacy class in $\mathcal{B}_{3}=\pi_{1}(\mathbb{U})$. A conjugacy class in the braid group on three strands defines, by taking its closure, a link $\sigma_{A}$ in a solid torus. In [Sim22a, Proposition 5.16] we relate the Alexander polynomial $\Delta(\sigma_{A})\in\mathbb{Z}[t^{\pm 1}]$ of this link $\sigma_{A}$ to the Fricke polynomial $\operatorname{Tr}A_{q}\in\mathbb{Z}[q^{\pm 1}]$ of the modular geodesic $[\alpha]$.
@@
The collection of $\operatorname{C}_{A}\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ for $A\in\mathcal{P}_{+}$ is linearly independant and every element $f\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ can be written as $f=\sum_{A\in\mathcal{P}_{+}}c_{A}(f).\operatorname{C}_{A}$ for unique $c_{A}(f)\in\mathbb{R}$.
To prove the non-triviality and linear independance of the $\operatorname{C}_{A}$ for $A\in\mathcal{P}_{+}$, we were led to show the non-degeneracy of the linking form, which is interesting in its own right.
@@ 1. The group $\operatorname{PSL}_{2}(\mathbb{R})$: discriminant and cross-ratio
The group $\operatorname{PSL}_{2}$ acts by conjugacy over itself. In this section, we recall from [Sim22b] the main invariants which enable to describe the orbits of single elements and of pairs of elements. Those are the discriminant $\operatorname{disc}(A)$ and the cross-ratio $\operatorname{bir}(A,B)$.
@@ 1.1. Over a field of characteristic $\neq 2$
We may also define their cosine (using their adjoint action on the Lie algebra $\operatorname{\mathfrak{sl}}_{2}(\mathbb{K})$ as in [Sim22b]), which is related to the cross-ratio by:
@@ 1.2. Over the real field
A hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{R})$ acts on $\mathbb{H}\mathbb{P}$ by translation along an (oriented) geodesic $\alpha$ whose endpoints $\alpha_{-},\alpha_{+}\in\mathbb{R}\mathbb{P}^{1}$ are its repulsive and attractive fixed points. With this order the period satisfies $\epsilon_{A}^{2}>1$. The translation length $\lambda_{A}=\log(\epsilon_{A}^{2})$ yields $\operatorname{disc}(A)=\left(2\sinh\tfrac{1}{2}\lambda_{A}\right)^{2}$
@@
If $\alpha$ and $\beta$ intersect, then it is with an angle $\theta\in]-\pi,\pi[$ oriented from $\alpha$ to $\beta$ according to $\operatorname{sign}(\theta)=\operatorname{sign}(A,B)$, and satisfying $\cos(\theta)=\cos(A,B)$, thus:
@@
$\tfrac{1}{\operatorname{bir}(\alpha,\beta)}=\tfrac{1+\cosh(\lambda)}{2}=\left(\cosh\tfrac{\lambda}{2}\right)^{2}\quad\mathrm{and}\quad\tfrac{1}{\operatorname{bir}(\alpha,\beta)}=\tfrac{1-\cosh(\lambda)}{2}=\left(\sinh\tfrac{\lambda}{2}\right)^{2}$
@@ 2.1. The modular orbifold
It is a hyperbolic two-dimensional orbifold, with conical singularities of order $2$ $3$ associated to the fixed points $i$ $j$ of $S$ $T$, and a cusp associated to the fixed point $\infty\in\partial\mathbb{H}\mathbb{P}$ of $R=TS^{-1}$.
@@
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ is the orbifold fundamental group of $\mathbb{M}$, so its conjugacy classes correspond to the free homotopy classes of (oriented) loops in $\mathbb{M}$. The elliptic conjugacy classes are those of $S$ $T^{\pm 1}$ which correspond to loops encircling the singularities, and the parabolic conjugacy classes are those of $R^{n}$ which correspond to loops encircling the cusp. The conjugacy class of a hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ corresponds to the free homotopy class of a unique closed geodesic $[\alpha]\subset\mathbb{M}$, and its length equals $\lambda_{A}=\log(\epsilon_{A}^{2})$. These are the so called modular geodesics.
@@ 2.2. Acting on a trivalent tree
Let $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ be primitive of infinite order. It acts on $\mathcal{T}$ by translation along an oriented geodesic $\alpha_{\mathcal{T}}$ called its combinatorial axis, with endpoints $\alpha_{-},\alpha_{+}\in\partial\mathcal{T}=\mathbb{R}\mathbb{P}^{1}$. Observe that $\alpha_{\mathcal{T}}$ passes through the oriented base edge $\vec{e}_{i}$ of $\mathcal{T}$ exactly when its endpoints satisfy $\alpha_{-}\leq 0\leq\alpha_{+}$, which is equivalent to saying that $A$ maps the base triangle $(0,1,\infty)$ to a triangle of the form $(\tfrac{b}{d},\tfrac{a+b}{c+d},\tfrac{a}{c})$ with $a,b,c,d\in\mathbb{N}$, in other terms that $A$ belongs to the monoid $\operatorname{PSL}_{2}(\mathbb{N})$ freely generated by $L\&R$. In that case, $\alpha_{\mathcal{T}}$ follows a periodic sequence of left and right turns given by the $L\&R$-factorisation of $A$, or the continued fraction expansion of the periodic number $\alpha_{+}$.
The conjugacy class of $A$ corresponds to the orbit of $\alpha_{\mathcal{T}}$ under the action of $\operatorname{PSL}_{2}(\mathbb{Z})$ on $\mathcal{T}$. Hence the conjugacy classes of non-elliptic elements in $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the cyclic words over the alphabet $\{L,R\}$, and the hyperbolic classes yield the cycles in which both letters appear. The linear representatives of such an $L\&R$-cycle parametrize the intersection of the corresponding conjugacy class with $\operatorname{PSL}_{2}(\mathbb{N})$, whose elements are called its Euclidean representatives. For infinite order $A\in\operatorname{PSL}_{2}(\mathbb{Z})$, the minimum displacement length $d(x,A\cdot x)$ of a vertex $x\in\mathcal{T}$ equals the combinatorial length $\operatorname{len}(A)=\#R+\#L$ of a Euclidean representative.
@@
In [Sim22a, Chapter 3], we recover the isotopy class of $[\alpha]\subset\mathbb{M}$ from the $L\&R$-cycle of $A$. We also describe when $[\alpha]$ passes through the singular points $i$ or $j$. Moreover [Sim22a, Lemma 2.27] shows that if a (primitive) hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})$ is stable under inversion, then it contains (exactly $4$) symmetric matrices, and those all lie in $\operatorname{PSL}_{2}(\mathbb{N})$ up to inversion.

### 04 Mar 2024 v4

@@ Abstract.
The linking numbers between modular knots and the trefoil are well understood. Indeed, É. Ghys showed in 2006 that they are given by the Rademacher invariant of the corresponding conjugacy classes. The Rademacher function $\operatorname{Rad}\colon\operatorname{PSL}_{2}(\mathbb{Z})\to\mathbb{Z}$ is a homogeneous quasi-morphism which he recognised with J. Barge in 1992 as half the primitive of the bounded Euler class.
@@
This paper contains the main results obtained in the second part of my thesis. I would thus like to thank my thesis advisors Etienne Ghys and Patrick Popescu-Pampu for their guidance and encouragement; as well as Francis Bonahon, Louis Funar, Jean-Pierre Otal and Anne Pichon who refereed and carefully read my work. I am also grateful to Pierre Dehornoy for sharing his knowledge of modular knots, as well as to the members of the arithmetics and dynamics teams at PSU for inviting me to present my works in their seminars. I owe Marie Dossin for helping me with the figures in tikz. Finally, I thank the referee for their relevant suggestions and questions.
@@ 0. Introduction
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ acts properly discontinuously on the hyperbolic plane $\mathbb{H}\mathbb{P}$ with quotient the modular orbifold $\mathbb{M}$, a hyperbolic surface with conical singularities $i$ and $j$ of order $2$ and $3$, and a cusp $\infty$. The free homotopy classes of loops in $\mathbb{M}$ correspond to the conjugacy classes of its fundamental group $\pi_{1}(\mathbb{M})=\operatorname{PSL}_{2}(\mathbb{Z})$. In particular the hyperbolic conjugacy classes of $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the closed oriented geodesics in $\mathbb{M}$, called modular geodesics. For hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ the modular geodesic $\alpha$ has length $\lambda_{A}$ equal to the logarithm of the ratio between its eigenvalues $\epsilon_{A}^{\pm 1}$, in formula:
@@
$L=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\qquad\mathrm{and}\qquad R=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$
and more precisely that its submonoid $\operatorname{SL}_{2}(\mathbb{N})$ of matrices with non-negative entries is freely generated by $\{L,R\}$. This submonoid can be identified with its image $\operatorname{PSL}_{2}(\mathbb{N})\subset\operatorname{PSL}_{2}(\mathbb{Z})$. In $\operatorname{PSL}_{2}(\mathbb{Z})$, the conjugacy class of an infinite order element intersects $\operatorname{PSL}_{2}(\mathbb{N})$ along all cyclic permutations of a non-empty $\{L,R\}$-word. The conjugacy class is primitive if and only if the cyclic word is primitive, and it is hyperbolic when the cyclic word contains both letters $L$ and $R$.
One may try to relate the geometry and topology of the master modular link with the arithmetics and combinatorics of conjugacy classes in the modular group. Our previous work [Sim23] relates the geometry of modular geodesics (angles of intersection and lengths of ortho-geodesics) in terms of the arithmetics of conjugacy classes in the modular group (discriminants, cross-ratios between fixed points on the projective line, and their Hilbert symbols). The main results in this paper will relate the linking numbers of modular knots to the combinatorics of the corresponding cyclic words.
@@
The arithmetic and geometry of the cosines
The journey from arithmetics began with our previous work [Sim23]. In particular for a field $\mathbb{K}\supset\mathbb{Q}$, we described when two hyperbolic elements $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ of the same discriminant $\Delta$ are conjugate in $\operatorname{PSL}_{2}(\mathbb{K})$. The obstruction is measured in terms of any intersection angle $\theta$ between their modular geodesics $[\alpha],[\beta]$ by the class of $\left(\cos\tfrac{\theta}{2}\right)^{2}\in\mathbb{K}^{\times}$ modulo the group of norms over $\mathbb{K}$ of the extension $\mathbb{K}[\sqrt{\Delta}]$. When this obstruction vanishes, the elements of $\operatorname{PSL}_{2}(\mathbb{K})$ which conjugate $A$ to $B$ are paremetrized by the points $(X,Y)\in\mathbb{K}$ of the generalised Pell-Fermat conic with equation $X^{2}-\Delta Y^{2}=\left(\cos\tfrac{\theta}{2}\right)^{2}$. Hence the geometric quantities $\left(\cos\tfrac{\theta}{2}\right)^{2}=\tfrac{1+\cos(\theta)}{2}$ given for some representatives $A,B\in\operatorname{SL}_{2}(\mathbb{R})$ whose axes intersect in $\mathbb{H}\mathbb{P}$ by
@@
$\operatorname{C}_{q}(A,B)\xrightarrow[q\to\infty]{}\operatorname{lk}(A,B)-\operatorname{lk}(A^{-1},B)=2\operatorname{lk}(A,B)-\tfrac{1}{2}I(A,B)$
Hence the functions $\operatorname{L}_{q}$ and $\operatorname{C}_{q}$ interpolate between the geometry at $q=1$ of the arithmetic group $\operatorname{PSL}_{2}(\mathbb{Z})\subset\operatorname{PSL}_{2}(\mathbb{R})$ and the topology at $q=\infty$ of the combinatorial action $\operatorname{PSL}_{2}(\mathbb{Z})\to\operatorname{Aut}(\mathcal{T})$.
@@
A primitive hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})=\pi_{1}(\mathbb{M})$ corresponds to a primitive modular geodesic in $\alpha\subset\mathbb{M}$. It lifts to a modular knot in $\vec{\alpha}\subset\mathbb{U}$ which in turns yields a conjugacy class in $\mathcal{B}_{3}=\pi_{1}(\mathbb{U})$. A conjugacy class in the braid group on three strands defines, by taking its closure, a link $\sigma_{A}$ in a solid torus. In [Sim22, Proposition 5.16] we relate the Alexander polynomial $\Delta(\sigma_{A})\in\mathbb{Z}[t^{\pm 1}]$ of this link $\sigma_{A}$ to the Fricke polynomial $\operatorname{Tr}A_{q}\in\mathbb{Z}[q^{\pm 1}]$ of the modular geodesic $[\alpha]$.
@@
The collection of $\operatorname{C}_{A}\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ for $A\in\mathcal{P}_{+}$ is linearly independent and every element $f\in PX(\operatorname{PSL}_{2}(\mathbb{Z}))$ can be written as $f=\sum_{A\in\mathcal{P}_{+}}c_{A}(f).\operatorname{C}_{A}$ for unique $c_{A}(f)\in\mathbb{R}$.
To prove the non-triviality and linear independence of the $\operatorname{C}_{A}$ for $A\in\mathcal{P}_{+}$, we were led to show the non-degeneracy of the linking form, which is interesting in its own right.
@@ 1. The group $\operatorname{PSL}_{2}(\mathbb{R})$: discriminant and cross-ratio
The group $\operatorname{PSL}_{2}$ acts by conjugacy over itself. In this section, we recall from [Sim23] the main invariants which enable to describe the orbits of single elements and of pairs of elements. Those are the discriminant $\operatorname{disc}(A)$ and the cross-ratio $\operatorname{bir}(A,B)$.
@@ 1.1. Over a field of characteristic $\neq 2$
We may also define their cosine (using their adjoint action on the Lie algebra $\operatorname{\mathfrak{sl}}_{2}(\mathbb{K})$ as in [Sim23]), which is related to the cross-ratio by:
@@ 1.2. Over the real field
A hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{R})$ acts on $\mathbb{H}\mathbb{P}$ by translation along an (oriented) geodesic $\alpha$ whose endpoints $\alpha_{-},\alpha_{+}\in\mathbb{R}\mathbb{P}^{1}$ are its repulsive and attractive fixed points. With this order the period satisfies $\epsilon_{A}^{2}>1$. The translation length $\lambda_{A}=\log(\epsilon_{A}^{2})$ yields $\operatorname{disc}(A)=\left(2\sinh\tfrac{1}{2}\lambda_{A}\right)^{2}$.
@@
If $\alpha$ and $\beta$ intersect, then it is with an angle $\theta\in]-\pi,\pi[$ oriented anti-clockwise from $\alpha$ to $\beta$ according to $\operatorname{sign}(\theta)=:\operatorname{sign}(A,B)$, and satisfying $\cos(\theta)=\cos(A,B)$, thus:
@@
$\tfrac{1}{\operatorname{bir}(A,B)}=\tfrac{1+\cosh(\lambda)}{2}=\left(\cosh\tfrac{\lambda}{2}\right)^{2}\quad\mathrm{and}\quad\tfrac{1}{\operatorname{bir}(A,B)}=\tfrac{1-\cosh(\lambda)}{2}=-\left(\sinh\tfrac{\lambda}{2}\right)^{2}$
@@ 2.1. The modular orbifold
It is a hyperbolic two-dimensional orbifold, with conical singularities of order $2$ and $3$ associated to the fixed points $i$ and $j$ of $S$ and $T$, and a cusp associated to the fixed point $\infty\in\partial\mathbb{H}\mathbb{P}$ of $R=TS^{-1}$.
@@
The modular group $\operatorname{PSL}_{2}(\mathbb{Z})$ is the orbifold fundamental group of $\mathbb{M}$, so its conjugacy classes correspond to the free homotopy classes of (oriented) loops in $\mathbb{M}$. The elliptic conjugacy classes are those of $S$ and $T^{\pm 1}$ which correspond to loops encircling the singularities, and the parabolic conjugacy classes are those of $R^{n}$ which correspond to loops encircling the cusp. The conjugacy class of a hyperbolic $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ corresponds to the free homotopy class of a unique closed geodesic $[\alpha]\subset\mathbb{M}$, and its length equals $\lambda_{A}=\log(\epsilon_{A}^{2})$. These are the so called modular geodesics.
@@ 2.2. Acting on a trivalent tree
Let $A\in\operatorname{PSL}_{2}(\mathbb{Z})$ be primitive of infinite order. It acts on $\mathcal{T}$ by translation along an oriented geodesic $\alpha_{\mathcal{T}}$ called its combinatorial axis, with endpoints $\alpha_{-},\alpha_{+}\in\partial\mathcal{T}=\mathbb{R}\mathbb{P}^{1}$. Observe that $\alpha_{\mathcal{T}}$ passes through the oriented base edge $\vec{e}_{i}$ of $\mathcal{T}$ exactly when its endpoints satisfy $\alpha_{-}\leq 0\leq\alpha_{+}$, which is equivalent to saying that $A$ maps the base triangle $(0,1,\infty)$ to a triangle of the form $(\tfrac{b}{d},\tfrac{a+b}{c+d},\tfrac{a}{c})$ with $a,b,c,d\in\mathbb{N}$, in other terms that $A$ belongs to the monoid $\operatorname{PSL}_{2}(\mathbb{N})$ freely generated by $\{L,R\}$. In that case, $\alpha_{\mathcal{T}}$ follows a periodic sequence of left and right turns given by the $\{L,R\}$-factorisation of $A$, or the continued fraction expansion of the periodic number $\alpha_{+}$.
The conjugacy class of $A$ corresponds to the orbit of $\alpha_{\mathcal{T}}$ under the action of $\operatorname{PSL}_{2}(\mathbb{Z})$ on $\mathcal{T}$. Hence the conjugacy classes of non-elliptic elements in $\operatorname{PSL}_{2}(\mathbb{Z})$ correspond to the cyclic words over the alphabet $\{L,R\}$, and the hyperbolic classes yield the cycles in which both letters appear. The linear representatives of such an $\{L,R\}$-cycle parametrize the intersection of the corresponding conjugacy class with $\operatorname{PSL}_{2}(\mathbb{N})$, whose elements are called its Euclidean representatives. For infinite order $A\in\operatorname{PSL}_{2}(\mathbb{Z})$, the minimum displacement length $d(x,A\cdot x)$ of a vertex $x\in\mathcal{T}$ equals the combinatorial length $\operatorname{len}(A)=\#R+\#L$ of a Euclidean representative.
@@
In [Sim22, Chapter 3], we recover the isotopy class of $[\alpha]\subset\mathbb{M}$ from the $\{L,R\}$-cycle of $A$. We also describe when $[\alpha]$ passes through the singular points $i$ or $j$. Moreover [Sim22, Lemma 2.27] shows that if a (primitive) hyperbolic conjugacy class in $\operatorname{PSL}_{2}(\mathbb{Z})$ is stable under inversion, then it contains (exactly $4$) symmetric matrices, and those all lie in $\operatorname{PSL}_{2}(\mathbb{N})$ up to inversion.